It took a little while to get going today.
  1. The Bayside X stage wouldn't home because we were not at 341, due to Dave's work. I had him move it to 341 and the stage homed.
  2. The processes disconnected from the message demon, or something. w_check showed them as being connected, but I got an error message saying that we had lost connection with the slope computer and all of the status icons on the HW GUI went red. I restarted the WFS and was able to home the stages.
  3. We couldn't find the light. The tertiary was out of the beam and then, later, I saw that the shutter on the ARGOS cal unit was closed.
20:48
  • We have the light.
  • Hexapod 2.05, 2.63, -1.5, -51.8, -209.49
  • Running optimize gain. The plots don't show on the FLAO GUI. Possibly because we haven't done a proper preset.
Here is the list of the reconstructors and their operational parameters, which we are going to test. Dave should be finished today, so hopefully we can do the proper procedure tomorrow.

/home/AOeng/aoroot/calib/wfs/current/ao/table_LUCIFER_ACE-AO.txt
# Modified Jun 06,2019 by SOUL team,    bin1,2 and 4 pupil changed shifting the center by Y +5, +3 and +1 pic respectivelydified Jun 06,2019 by SOUL team,    bin1,2 and 4 pupil changed shifting the center by Y +5, +3 and +1 pic respectively
# Mofidied Jun 07,2019 by SOUL team,    changed FW#2 for acq to 650-10
# Modified Jul 04,2019 by E. Pinna,     added 3 columns for Gopt modulation config (all at 30 10 80)
#                                       changed BIN4 REC and KL (KL20->KL18) to align to the one used for ARGOS (ROT offset 225 not 350)
# Modified Jul 09,2019 by E. Pinna,     changed gopt modulation amplitude for 14 to 15.5 up to 15 instead of 10nm
# Modified Oct 18,2019 by G. Agapito,   changed all filterings with new iir filters and forgetting factors, and some rec matrix to be compatible with iir filters
# Modified Oct 23,2019 by G. Agapito,   iir filter are now old ones already tested
# Modified Oct 24,2019 by G. Agapito,   added few lines to manage correctly transition between IIR filters (R=11.4, 11.7, 12.5, 12.7, 14.7, 15.0, 15.2)
# Modified Nov 07,2019 by F. Rossi,     changed version of IIR filter at 11.4
# Modified Dec 13,2019 by G. Agapito,    changed bin 1 rec Rmag>12.5 (300 modes), bin 4 rec and pupil files Rmag>=15.2
#                     changed filtering Rmag>=15.2, now it is forgetting factors
#                     modulation signal amp. for bin 2 @Rmag=13.0 and all mags for bin 4
# Modified Feb 07,2020 by E. Pinna       New bin4 pupil shifted (0.3pix) changed dist amp to 30e-9 for 14.5 both bin2 and 4
# Modified Feb 27,2020 by E. Pinna      Bin3 implemeted following test of 26 and 27 Feb 2020
# Modified July 2,2020 by E. Pinna     BIN1 -> BIN2 moved form 13.0 to 12.7; the bin111 300 modes 500Hz was not smart WRT bin2 250 modes 1kHz
#
# Columns:
# 0    1      2   3     4      5      6                7                          8                 9    
Rmag   flux   bin freq  emGain base   rec              filtering                  pupils            tt   

-1.5 1.054E9  1  1700  1       KL_v20 20190930_161000  iir1700HzV003              20190909_203854   3    
0.0   2.65E8  1  1700  1       KL_v20 20190930_161000  iir1700HzV003              20190909_203854   3    
3.5  10.54E6  1  1700  10      KL_v20 20190930_161000  iir1700HzV003              20190909_203854   3    
4.5   4.20E6  1  1700  20      KL_v20 20190930_161000  iir1700HzV003              20190909_203854   3    
5.5   1.67E6  1  1700  30      KL_v20 20190930_161000  iir1700HzV003              20190909_203854   3    
6.5   6.65E5  1  1700  100     KL_v20 20190930_161000  iir1700HzV003              20190909_203854   3    
7.5   2.65E5  1  1700  100     KL_v20 20190930_161000  iir1700HzV003              20190909_203854   3    
8.5   1.05E5  1  1700  100     KL_v20 20190930_161000  iir1700HzV003              20190909_203854   3    
9.5   4.20E4  1  1700  300     KL_v20 20190930_161000  iir1700HzV003              20190909_203854   3    
10.5  1.67E4  1  1700  600     KL_v20 20190930_161000  iir1700HzV003              20190909_203854   3    
11.4  7.30E3  1  1350  600     KL_v20 20190930_161000  iir1700HzV003              20190909_203854   3    
11.4  7.30E3  1  1350  600     KL_v20 20190930_161000  iir1350HzV001              20190909_203854   3    
11.7  5.53E3  1  1000  600     KL_v20 20190930_161000  iir1350HzV001              20190909_203854   3    
11.7  5.53E3  1  1000  600     KL_v20 20190930_161000  iir1000HzV001              20190909_203854   3    
12.5    2648  1   750  600     KL_v20 20190930_161000  iir1000HzV001              20190909_203854   3    
12.5    2648  1   750  600     KL_v20 20191213_000000  iir0750HzBin1              20190909_203854   3    
12.7    2202  1   625  600     KL_v20 20191213_000000  iir0750HzBin1              20190909_203854   3    
12.7    2202  2  1250  600     KL_v20 20191011_125000  iir1350HzV001              20190606_074345   3    
13.0    1671  2  1000  600     KL_v20 20191011_125000  iir1000HzV001              20190606_074345   3    
14.0     665  2   800  600     KL_v20 20191011_125000  iir1000HzV001              20190606_074345   3    
14.5     420  2   750  600     KL_v20 20191011_125000  iir1000HzV001              20190606_074345   3    
14.5     420  3  1000  600     KL_v20 20200226_000000  iir1000HzV001              20200210_000000   3    
15.5     167  3   750  600     KL_v20 20200226_000000  iir1000HzV001              20200210_000000   3    
15.5     167  3   750  600     KL_v20 20200226_000000  iir0750HzV001              20200210_000000   3    
15.8     135  3   500  600     KL_v20 20200226_000000  iir0750HzV001              20200210_000000   3    
15.8     135  3   500  600     KL_v20 20200225_144315  forgettingFactorBin4SOUL   20200210_000000   3    
16.5      67  3   250  600     KL_v20 20200225_144315  forgettingFactorBin4SOUL   20200210_000000   3    
16.5      67  4   250  600     KL_v18 20200224_223616  forgettingFactorBin4SOUL   20200113_010101   3    
17.5    26.5  4   100  600     KL_v18 20191209_201936  forgettingFactorBin4SOUL   20200113_010101   3   
18.5    10.5  4   100  600     KL_v18 20191209_201936  forgettingFactorBin4SOUL   20200113_010101   3   
19.5     4.2  4   100  600     KL_v18 20191209_201936  forgettingFactorBin4SOUL   20200113_010101   3 
20.5     1.7  4   100  600     KL_v18 20191209_201936  forgettingFactorBin4SOUL   20200113_010101   3   

Results for each Reconstructor:

We had a problem with the 2" distrubance file dist_flao2_KL_v20_atm_s2.0_L040.0_v15.0_ovfreq1700.00_sd3892.fits. It causes forces that are too high and the loop won't stay closed. Guido suggests that this is a problem with the file itself and not with the inherent amplitude of 2" seeing.

Binning KL Basis Reconstructor Frequency iir Mode em Gain Measured Magnitude ARGOS Shutter Disturbance Pass/Fail Tracking Number
1 KL_v20 20190930_161000 1700 iir1700HzV003 600 10.6 303 (fully Open)

/local/aomeas/adsec_calib/CMD/disturb/

dist_flao2_KL_v20_atm_s1.0_L040.0_v15.0_ovfreq1700.00_sd3892.fits

Pass
20200828_223206
1 KL_v20 20190930_161000 1350 iir1350HzV001 600
1 KL_v20 20190930_161000 1000 iir1000HzV001 600
1 KL_v20 20191213_000000 750 iir0750HzBin1 600
2 KL_v20 20191011_125000 1250 iir1350HzV001 600
2 KL_v20 20191011_125000 1000 iir1000HzV001 600
3 KL_v20 20200226_000000 1000 iir1000HzV001 600
3 KL_v20 20200226_000000 750 iir0750HzV001 600
3 KL_v20 20200225_144315 500 forgettingFactorBin4SOUL 600
4 KL_v18 20200224_223616 250 forgettingFactorBin4SOUL 600
4 KL_v18 20191209_201936 100 forgettingFactorBin4SOUL 600
-- GregoryTaylor - 28 Aug 2020
Topic revision: r2 - 29 Aug 2020, GregoryTaylor
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback